OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 2x - y = 8. intercept = (2, 0); y-intercept = (0, -8) x-intercept = (4, 0); y-intercept = (0, -8) intercept = (2, 0); y-intercept = (0, 8) intercept = (4, 0); y-intercept = (0, 8)

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

2x-y=8 x/4-y/8=1 x/4+y/-8=1 x-intercept = (4, 0); y-intercept = (0, -8)

OpenStudy (anonymous):

yea, for x-intercept put y=0 and y-intercept put x=0